# GEO: Generative Engine Optimization

### v2

@@ 8. Conclusion
In this work, we formulate search engines augmented with generative models that we dub generative engines. We propose Generative Engine Optimization (GEO) to empower content creators to optimize their content under generative engines. We define impression metrics for generative engines and propose and release GEO-bench: a benchmark encompassing diverse user queries from multiple domains and settings, along with relevant sources needed to answer those queries. We propose several ways to optimize content for generative engines and demonstrate that these methods are capable of boosting source visibility by up to 40% in generative engine responses. Among other findings, we show that including citations, quotations from relevant sources, and statistics can significantly boost source visibility. Further, we discover a dependence of the effectiveness of GEO methods on the domain of the query, and potential of combining multiple GEO strategies in conjunction. We show promising results on a commercially deployed generative engine with millions of active users showcasing the real-world impact of our work. Our work serves as a first step towards understanding the impact of generative engines on the digital space and the role of GEO in this new paradigm of search engines.
@@ 9. Limitations
While we rigorously test our proposed methods on two generative engines, including a publicly available one, methods may need to adapt over time as GEs evolve, mirroring the evolution of SEO. Additionally, despite our efforts to ensure the queries in our GEO-bench closely resemble real-world queries, the nature of queries can change over time, necessitating continuous updates. Lastly, while every query in our proposed GEO-bench is tagged and manually inspected, there may be discrepancies due to subjective interpretations or errors in labeling.
@@ Appendix A Conversational Generative Engine
Further, to engage the user in a conversation, a separate LLM, or , may generate suggested follow-up queries based on , , and . The suggested follow-up queries are typically designed to maximize the likelihood of user engagement. This not only benefits Generative Engine providers by increasing user interaction, but also benefits website owners by enhancing their visibility. Furthermore, these follow-up queries can help users by getting more detailed information.
@@ B.1. Evaluated Generative Engine
While the exact design specifications of popular Generative Engines are not public, based on previous works (Liu et al., 2023b), and our experiments, most of them follow 2-step procedure for generating responses. The first step involves fetching relevant sources for input query, followed by a LLM generating response based on the fetched sources. In our setup, we fetch top 5 sources from Google search engine for every query. The answer is generated by gpt3.5-turbo model using a prompt same as prior work (Liu et al., 2023b). The model is prompted to output an appropriate response for the query, with each sentence cited by one of the 5 sources provided. We sample 5 different answers at temperature=0.7 and top_p=1, to reduce statistical deviations. The exact prompt used is shown in Listing 1.
@@ B.2. Benchmark
Since, currently there is no publicly available dataset containing Generative Engine related queries we curate GEO-bench, a benchmark containing 10K queries from multiple sources repurposed for generative engines along with synthetically generated queries. The benchmark contains queries from nine different sources, each further categorized based on their target domain, difficulty, query intent and other dimensions. The datasets used in constructing the benchmark are: 1. MS Macro & 2. ORCAS-1: contains real anonymized user queries from Bing Search Engine. and 3. Natural Questions: containing queries from Google Search Engine. These three collectively represent the common set of datasets that are used in search engine related research. However, Generative Engines will be posed with far more dofficult and specific queries with intent of synthesizing answer from multiple sources instead of search for them. To this end, we re-purpose several other publicly available datasets: 4. AllSouls: A dataset containing essay questions from “All Souls College, Oxford University”. The queries in the dataset cannot be usually answered from a single source, and requires Generative Engines to aggregate information multiple sources and perform reasonable reasoning on them. 5. LIMA (Zhou et al., 2023) contains carefully crafted queries and responses for training pretrained language models for instruction following. The queries in the dataset represent a more challenging distribution of queries asked in Generative Engine, and often requires LLM’s creative and technical powress to generate answers (eg: writing short poem, python code.) 6. Perplexity.ai Discover: These queries are sourced from Perplexity.ai’s, a public Generative Engine, Discover section which is a updated list of trending queries on the platform. These queries represent a real distribution of queries made on Generative Engines. 7. Davinci-Debtate (Liu et al., 2023b) contains debate questions generated using text-davinci-003 and sourced from Perspectrum dataset (Chen et al., 2019). This dataset were specifically designed for Generative Engines. 8. ELI-5: contains questions from the ELI5 subreddit, where users ask complex questions and expect answers in simple, layman terms. 9. GPT-4 Generated Queries: To further supplement diversity in query distribution and increase Generative Engine specific queries, we prompt GPT-4 to generate queries ranging from various domains (eg: science, history), based on query intent (eg: navigational, transactional), based on difficulty and scope of generated response (eg: open-ended, fact-based). Representative queries from each of the datasets are shown in Figure 2.
In total our benchmark contains 10K queries split into 8K,1K,1K train/val/test splits. Every query is tagged into multiple categories gauging various dimensions such as intent, difficulty, domain of query and format of answer type using GPT-4. In terms of query intent, we maintain the real-world query distribution, with our benchmark containing 80% informational queries, and 10% transactional queries and 10% navigational queries (Jansen et al., 2008). Further, we augment every query with cleaned text content of top 5 search results from Google search engine. Owing to specially designed high benchmark diversity, size, complexity and real-world nature, we believe GEO-bench is a comprehensive benchmark for evaluating Generative Engines and serves as a standard testbed for evaluating Generative Engines for multiple purposes in this and future works.
Tags:
Optimizing website content often requires making targeted changes based on the domain of the task. Further, a user of Generative Engine Optimization may need to find an appropriate method for only a subset of queries based on multiple factors, such as domain, user intent, query nature. To this end, we tag each of the queries based from a pool of 7 different categories. For tagging we use GPT-4 model, and manually confirm high recall and precision in tagging. However, owing to such automated system, the tags can be noisy and should not be considered. Detail about each of these queries are presented here and Listing 5
@@
1# Example of a simple query2query = "What is the capital of France?"3# Example of a complex query4query = "What are the implications of the Schrodinger equation in quantum mechanics?" -
@@
1# Example of a factual query2query = "How does a car engine work?"3# Example of an opinion query4query = "What is your opinion on the Harry Potter series?" -
@@
1# Example of a query in the arts and entertainment genre2query = "Who won the Oscar for Best Picture in 2020?"3# Example of a query in the finance genre4query = "What is the current exchange rate between the Euro and the US Dollar?" -
@@
1# Example of a query on a specific topic in physics2query = "What is the theory of relativity?"3# Example of a query on a specific topic in economics4query = "What is the law of supply and demand?" -
@@
1# Example of a non-sensitive query2query = "What is the tallest mountain in the world?"3# Example of a sensitive query4query = "What is the current political situation in North Korea?" -
@@
1# Example of a research intent query2query = "What are the health benefits of a vegetarian diet?"3# Example of a purchase intent query4query = "Where can I buy the latest iPhone?" -
@@
1# Example of a fact answer type query2query = "What is the population of New York City?"3# Example of an opinion answer type query4query = "Is it better to buy or rent a house?"
@@ B.3. Evaluation Metrics
We evaluate all methods by measuring the Relative Improvement in Impression. Specifically, given an initial generated Response from sources s, and a modified response from sources s, we measure the relative improvement in impression of each of the source as:
We use the impression metrics as defined in Section 2.2.1. Specifically, we use two impression metrics: 1. Position-Adjusted Word Count which is a combination of word count and position count. To dissect effect of individual components, we also report individual scores on the 2 sub-metrics. 2. Subjective Impression which is a subjective impression metric which is a combination of seven different aspects: Relevance of citation to query, influence of citaiton on response, diversity and uniqueness of information presented, likelihood of followup by the user, perceived rank and amount of information presented in the answer. All these sub-metrics are evaluated using GPT-3.5, using methodology similar to described as in G-Eval (Liu et al., 2023a). However, since G-Eval scores are ill-calibrated, we need to suitably normalize them for fair and appropriate comparison. We normalize Subjective Impression scores with respect to baseline scores of Position-Adjusted Word Count to ensure same mean and standard deviation. We list prompts for each of the sub-metrics in Listing 6 7 8 9.
B.4. GEO Methods
To improve the impression metrics, content creators need to make changes to their websites. To this end, we present several generative engine-agnostic strategies, refered to as Generative Engine Optimization methods (GEO). Mathematically, every GEO method is a function , where is the initial web content, and is the modified website content after applying LEO method. We propose and evaluate a series of methods.
1: Authoritative: Modifies text style of the source content to be more persuasive while making authoritative claims, 2. Statistics Addition: Modifies content to include quantitative statistics instead of qualitative discussion, wherever possible, 3. Keyword Stuffing: Modifies content to include more keywords from the query, as would be expected in classical SEO optimization. 4. Cite Sources & 5. Quotation Addition: Adds relevant citations and quotations from credible sources, 6.) 6. Easy-to-Understand: Simplifies the language of website, while 7. Fluency Optimization improves the fluency of website text. 8. Unique Words & 9. Technical Terms: involves adding unique and technical terms respectively wherever posssible,
In order to analyse the performance gain of our methods, for each input query, we randomly select one source to be optimized using all GEO separately. Further, for every method, 5 answers are generated per query to reduce statistical noise in the results.
B.5. Prompts for GEO methods
Here we present prompts used for different Generative Engine Optimization methods. GPT-3.5 turbo was used for all experiments.
Appendix C Results
We perform experiments on 5 seeds, and present results in Table 6
C.1. GEO in the Wild : Experiments with Deployed Generative Engine
To further reinforce the efficacy of our proposed Generative Engine Optimization methods, we evaluate them on Perplexity.ai a deployed Generative Engine with a large user base. Since perplexity.ai does not allow the user to specify source URLs, we instead provide source text as file uploads to perlexity.ai. We ensure all answers are generated only using the file sources provided. We evaluate all our methods on a subset of 200 samples of our test set. Other experimental procedures are the same as our main results. Results using Perplexity.ai are shown in Table 7. We find, similar to our generative engine Quotation Addition performs the best in Position-Adjusted Word Count with a relative improvement of 22% over the baseline. Further, methods that performed well in our generative engine such as Cite Sources, Statistics Addition show high improvements of up to 9% and 37% on the two metrics. Further, our observations such as the ineffectiveness of traditional methods used in SEO such as Keyword Stuffing are further highlighted, as it performs 10% worse than the baseline. The result underscores the importance of developing different Generative Engine Optimization methods to benefit the content-creators and further highlights that our simple-to-implement proposed methods can be used directly by content-creators, thus having a high real-world impact.
Appendix D Ethical Considerations and Reproducibility Statement
In our study, we focus on enhancing the visibility of websites in generative engines. We do not directly interact with sensitive data or individuals. While the sources we retrieve from search engines may contain biased or inappropriate content, these are already publicly accessible, and our study neither amplifies nor endorses such content. We believe that our work is ethically sound as it primarily deals with publicly available information and aims to improve the user experience in generative engines.
Regarding reproducibility, we have made our code available to allow others to replicate our results. Our main experiments have been conducted with five different seeds to minimize potential statistical deviations.

### v3

@@ 8. Conclusion
In this work, we formulate search engines augmented with generative models that we dub generative engines. We propose Generative Engine Optimization (GEO) to empower content creators to optimize their content under generative engines. We define impression metrics for generative engines and propose and release GEO-bench: a benchmark encompassing diverse user queries from multiple domains and settings, along with relevant sources needed to answer those queries. We propose several ways to optimize content for generative engines and demonstrate that these methods can boost source visibility by up to 40% in generative engine responses. Among other findings, we show that including citations, quotations from relevant sources, and statistics can significantly boost source visibility. Further, we discover a dependence of GEO methods’ effectiveness on the query domain and the potential of combining multiple GEO strategies in conjunction. We show promising results on a commercially deployed generative engine with millions of active users, showcasing the real-world impact of our work. In summary, our work is the first to formalize the important and timely GEO paradigm, releasing algorithms and infrastructure (benchmarks, datasets, and metrics) to facilitate rapid progress in generative engines by the community. This serves as a first step towards understanding the impact of generative engines on the digital space and the role of GEO in this new paradigm of search engines.
@@ 9. Limitations
While we rigorously test our proposed methods on two generative engines, including a publicly available one, methods may need to adapt over time as GEs evolve, mirroring the evolution of SEO. Additionally, despite our efforts to ensure the queries in our GEO-bench closely resemble real-world queries, the nature of queries can change over time, necessitating continuous updates. Further, owing to the black-box nature of search engine algorithms, we didn’t evaluate how GEO methods affect search rankings. However, we note that changes made by GEO methods are targeted changes in textual content, bearing some resemblance with SEO methods, while not affecting other metadata such as domain name, backlinks, etc, and thus, they are less likely to affect search engine rankings. Further, as larger context lengths in language models become economical, it is expected that future generative models will be able to ingest more sources, thus reducing the impact of search rankings. Lastly, while every query in our proposed GEO-bench is tagged and manually inspected, there may be discrepancies due to subjective interpretations or errors in labeling.
10. Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 2107048. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Conversational Generative Engine
Further, to engage the user in a conversation, a separate LLM, or , may generate suggested follow-up queries based on , , and . The suggested follow-up queries are typically designed to maximize the likelihood of user engagement. This not only benefits Generative Engine providers by increasing user interaction but also benefits website owners by enhancing their visibility. Furthermore, these follow-up queries can help users by getting more detailed information.
@@ B.1. Evaluated Generative Engine
The exact prompt used is shown in Listing 1.
@@ B.2. Benchmark
GEO-bench contains queries from nine datasets. Representative queries from each of the datasets are shown in Figure 2. Further, we tag each of the queries based on a pool of 7 different categories. For tagging, we use the GPT-4 model and manually confirm high recall and precision in tagging. However, owing to such an automated system, the tags can be noisy and should not be considered carefully. Details about each of these queries are presented here:
@@ B.3. Evaluation Metrics
We use 7 different subjective impression metrics, whose prompts are presented in our our public repository: https://github.com/GEO-optim/GEO.
@@
B.4. GEO Methods
We propose 9 different Generative Engine Optimization methods to optimize website content for generative engines. We evaluate these methods on the complete GEO-bench test split. Further, to reduce variance in results, we run our experiments on five different random seeds and report the average.
B.5. Prompts for GEO methods
We present all prompts in our our public repository: https://github.com/GEO-optim/GEO. GPT-3.5 turbo was used for all experiments.
Appendix C Results
We perform experiments on 5 random seeds and present results with statistical deviations in Table 6
C.1. GEO in the Wild : Experiments with Deployed Generative Engine
We also evaluate our proposed Generative Engine Optimization methods on real-world deployed Generative Engine: Perplexity.ai. Since perplexity.ai does not allow the user to specify source URLs, we instead provide source text as file uploads to perplexity.ai while ensuring all answers are generated only using the file sources provided. We evaluate all our methods on a subset of 200 samples of our test set. Results using Perplexity.ai are shown in Table 7.